OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point y=2x+4;(3,8) A; y=2x+2 B; y=2x+6 C; y=-2x+6 D; y=-1/2x+2

OpenStudy (ahsome):

Really simple :) When a line is parallel to another point, it basicly means it has the same slope Look at the line below:  A line that is parallel goes in the same direction, and therefore has the same slope 

OpenStudy (ahsome):

Now, you are aware of the basic equation of a line, aren't you? \(y=mx+c\) or \(y=mx+b\) Where: \(y\) Is the \(y\) value of a point \(x\) Is the \(x\) value of a point \(m\) Is the slope \(b\) Is a constant or number that is equal to the y-intercept Now, we know y and x, since you are given a point. We also know the slope, since you said that the line is parallel to y=2x+4. In this equation, the slope or \(m\) value is 2. Therefore, in our new equation, the slope or \(m\) value must be 2 also. Now since we know 3 out of 4 numbers, we can use algebra to find the last number \(b\) \[y=mx+n\]\[8=2\times3+b\]\[8=6+b\]\[8-6=b\]\[2=b\]\[b=2\] Sub the values we know \(m\) and \(b\) \[y=mx+b\]\[y=2x+2\] Thats the answer :)

OpenStudy (anonymous):

thank you, thank you

OpenStudy (ahsome):

No problem @penguinlol88 :)

OpenStudy (ahsome):

Do you have any more questions we could help you with, @penguinlol88?

OpenStudy (ahsome):

I'm taking that as a no ;) @penguinlol88

OpenStudy (anonymous):

Describe how the graphs of y=/x/ and y=/x+3/ are related. A; the two graphs are the same B; the graphs have the same shape the y-intercept of x=/x/is 0and the x-intercept of the second graph is 3 C; the graphs have the same shape the y-intercept of y=/x/ is 0 and the x-intercept of the second graph is 3 d; the graphs have the same shape the y-intercept of y=/x/is 0 and the y-intercept of the second graph is -3

OpenStudy (ahsome):

Ask this in a new question, then we all can help you :)